x+y+2=1
3x-2y+42=-19
2x+5y-32=2

The directions are
"Solve the system. Record the answer is the appropriate manner:"

Answer by ikleyn(52777)

.

It seems to me that the third variable/unknown " z " is missed in all three equations.

Double check with your source.


---------------

Comment from student: Sorry this is the problem is x+y+z=1 3x-2y+4z=-19 2x+5y-3z=2


My response : It is just much better.

Please re-post the correct problem to the forum.
Do not post it to me personally.


Actually, this windows "Comment from student" is intended for one single purpose: to express your "THANKS" to a tutor.

It is not intended to post new problems or new versions of the problems to tutors.


have a happy holiday season (!)